OpenStudy (anonymous):

HELP I ONLY NEED 4MORE MATH PROBLEMS ...!!!:?

OpenStudy (anonymous):

(3^SQRT3-2)(3^SQRT9+2 3^SQRT3+4)=

OpenStudy (anonymous):

Rationalize the denominator. SQRT X + SQRT Y over SQRT X - SQRT Y=

myininaya (myininaya):

\[\frac{\sqrt{x}+\sqrt{y}}{\sqrt{x}-\sqrt{y}}*\frac{\sqrt{x}+\sqrt{y}}{\sqrt{x}+\sqrt{y}}\] \[=\frac{x+2\sqrt{x}\sqrt{y}+y}{x-y}\]

OpenStudy (anonymous):

Perform the division. X^2-1 divided into x^5+2x^3-x^2-3x-6=